OpenStudy (anonymous):

input he equation of the gven line in standard form. The line with m= -1/2 and the point (-2,-3)

OpenStudy (anonymous):

The formula is, y - y1 = m*(x - x1) so, y - (-3) = -1/2*{x - (-2)} y + 3 = -1/2*(x + 2) 2y + 6 = -x -2 2y + x + 8 = 0...would be the equation

OpenStudy (anonymous):

did that help? and was that correct? can u verify the answer?